# LOB-Bench: Benchmarking Generative AI for Finance – an Application to Limit Order Book Data

### v1

@@ 6 Results
Figure 4: Model score summaries (lower is better). The LOBS5 model achieves the lowest overall scores. coletta beats the baseline on the Wasserstein metric, but not for L1. Error bars are bootstrapped 99% CIs.
As a first test case for our benchmark, we have adapted the autoregressive state-space model using S5 layers (Gu et al., 2021) from Nagy et al. (2023) (LOBS5). Particularly, we have scaled up the model size to 35 million parameters and more than doubled the training period to the entire year of 2022.
We also evaluated data generated by the models from Cont et al. (2010) (baseline), Coletta et al. (2022) (Coletta) and based on Peng et al. (2023, 2024) (RWKV 4 and 6). The baseline model, which employs parametric arrival processes, was adapted to generalize across both small and large tick limit order book (LOB) dynamics by utilizing estimated empirical arrival rates directly, rather than fitting a power law. Additionally, we inferred data features present in LOBSTER, such as individual message IDs, which were not generated by Cont et al. (2010). This inference is particularly important for capturing order cancellations, as we uniformly sample target limit orders from the available orders at the specified price level. For the Coletta model, we implemented a LOBSTER data interface to facilitate the conversion of data formats. For the RWKV models, we apply autoregressive next-token prediction, but on a larger model (170 million parameters), without any data pre-processing and using an off-the-shelf byte-pair tokenizer, as is used for LLMs. These models are trained solely on message data, without any order books, requiring no propagation of a calculated orderbook state unlike in Nagy et al. (2023). The S5, RWKV, and baseline results presented here were computed, following Nagy et al. (2023), for Alphabet (GOOG) and Intel (INTC) stock on a sub-sample of the test data from January 2023. The Coletta model was trained on three days from January 2019 and tested on three subsequent days, following the procedure in Coletta et al. (2022), necessitated by the high computational cost for training and inference with Coletta. Comparing all models, we note that the LOBS5 model provides state-of-the-art performance on the benchmark task.
Figure 3 presents a key benchmark feature to compare multiple models across multiple score dimensions, allowing an examination of individual strengths and weaknesses. To provide summary scores per model, Figure 4 reports the mean, median, and inter-quartile mean for the L1 and Wasserstein-1 metrics for all models  [^2] The Coletta model (Coletta et al., 2022) was trained on both GOOG and INTC data, but failed to produce reasonable results for INTC, which is explainable as it was intended for small-tick stocks, which INTC is not.. Error bars demarcate the 99% bootstrapped confidence intervals. Metrics for individual scoring functions are shown in Figure 10 (Appendix D).
The benchmark also measures error divergence by comparing distributions of scoring functions, conditional on the inference time step. These demonstrate the rate at which distributions diverge from real data. Results show increasing errors across all models with the fastest divergence exhibited by the RWKV models across most scores. For the S5 model in particular, scoring functions with a dependence on features of the book states, which only gradually change, such as book volume, are expected to diverge, as the initial real data seed decays. However, the rate of decay can still be compared between models. See Figure 16 in the appendix for L1 divergence curves.
@@
Figure 5: LOBS5 results - (left): histogram matching of unconditional score distributions for real and generated data. (right): Error accumulation - the further out the prediction horizon, the worse is the model performance - an important model characteristic to measure.
@@
Figure 6: L1 distance between real and generated data histograms (99% bootstrapped CIs in black). The baseline performs well on LOB depth and level-related scores, and much worse on time and volume metrics. LOBS5 dominates L1 loss for GOOG
@@
Figure 7: Comparison of impact response functions for different event types between real and generated data-sets, tick-normalized mid-price response. Shaded regions are 99% confidence intervals. There is a comparison between two select models: the LOBS5 and the stochastic baseline. We see that, in contrast to the baseline, the generative model is able to reproduce most of the expected impact response function.
@@
The response functions for Alphabet (GOOG) are shown in  Figure 7 for the Baseline and LOBS5 models. The LOBS5 model generally reproduces curves similar to real data but does so better for small-tick stock GOOG. In contrast, the baseline model (Cont et al., 2010) cannot faithfully reproduce impact curves. We do not post the impact curves for the Coletta and RWKV models, as they very strongly diverge, due to the error accumulation at inference. For instance, the average L1 distance between the real and generated curves (7) is 2.45 for LOBS5, and 126 for RWKV-6. For the LOBS5 model we observe differences largely in the $MO$ orders at short lags. This is due to the treatment of the JAX-LOB (Frey et al., 2023) simulator at inference time, as used by the LOBS5 model. This interprets some large messages as execution and an additional limit order, merging a multi-level mid-price change into a single order book update.
@@ 7 Conclusions
One of the interesting aspects of generative AI models for microstructure data is the ability to model counterfactuals, which closely relates to the notion of price impact in financial modelling. Factoring in the reactions of other market participants to one’s actions with conventional approaches is very challenging, but our benchmark suite for generative LOB models provides extensive tests to evaluate whether generated data reproduces the expected response functions at a larger scale. In the future we plan to measure the extent to which generative models match the market impact laws in literature, such as the “square root law” (SRL) (Tóth et al., 2016). We hope that this opens the door to many new studies, including training of reinforcement learning algorithms and multi-agent models for trade execution with the ability to model realistic reactions of different market participants.
@@ Acknowledgements
We gratefully acknowledge the Oxford-Man Institute of Quantitative Finance and the Foerster Lab for AI Research for providing us access to their GPU compute servers. These computational resources were instrumental in carrying out the experiments and analyses presented in this work. AC acknowledges funding from a UKRI AI World Leading Researcher Fellowship (grant EP/W002949/1. AC and JF acknowledge funding from a JPMC Faculty Research Award.
@@ Appendix A Benchmark Code
The benchmark code can be found on GitHub at https://github.com/peernagy/lob_bench.
The benchmark suite provides a convenient API functionality to evaluate model data for a range of scoring functions and metrics. A specification of such functions and loss metrics can be defined in a configuration dictionary, which can then be passed to function performing the unconditional and conditional model evaluation. Similarly, the benchmark provides functions to compute the market impact curves, along with a mean L1 score. A default configuration dictionary, specifying the scoring functions reported here, evaluated using L1 and Wasserstein-1 loss, is similarly provided for easy reproducibility.
@@ Appendix B LOBS5 Training Details
Starting from the model introduced by Nagy et al. (2023), we have scaled up the model size by adding additional S5 layers, with a resulting parameter count of approximately 35M (compared to originally 6.3M). The training consisted of 100 epochs of shuffled data sequences from the entire year of 2022, training with a total training budget of 30.4 L40 days (3.8 days across 8 GPUs). Adam (Kingma & Ba, 2014) was used as an optimizer with a cosine learning rate schedule.
@@ Appendix C RWKV Training Details
Appendix D Additional Figures
@@
Figure 13: rwkv4 - histograms comparing score distributions for real (blue) and generated (orange) LOB data for Alphabet (GOOG) and Intel (INTC) stocks. The mode produces volatile data with larger spreads, missing correct order levels, leading to difficulty matching book volumes.

### v2

@@ 6 Results
Figure 5: LOBS5 results – (left): histogram matching of unconditional score distributions for real and generated data. (right): error accumulation – the further out the prediction horizon, the worse is the model performance – an important model characteristic to measure.
As a first test case for our benchmark, we adapt the autoregressive state-space model using S5 layers (Gu et al., 2021) from Nagy et al. (2023) (LOBS5). Particularly, we scale up the model size to 35 million parameters and more than double the training period to the entire year of 2022.
We also evaluate data generated by the models from Cont et al. (2010) (baseline), Coletta et al. (2022) (Coletta) and models based on Peng et al. (2023, 2024) (RWKV 4 and 6). The baseline model, which employs parametric arrival processes, is adapted to generalize across both small and large tick limit order book (LOB) dynamics by utilizing estimated empirical arrival rates directly, rather than fitting a power law. Additionally, we infer data features present in LOBSTER, such as individual message IDs, which are not generated by Cont et al. (2010). This inference is particularly important for capturing order cancellations, as we uniformly sample target limit orders from the available orders at the specified price level. For the Coletta model, we implement a LOBSTER data interface to facilitate the conversion of data formats. For the RWKV models, we apply autoregressive next-token prediction, but on a larger model (170 million parameters), without any data pre-processing and using an off-the-shelf byte-pair tokenizer (Sennrich et al., 2016), as is used for LLMs. These models are trained solely on message data, without any order books, which do not require propagation of a calculated orderbook state, unlike in Nagy et al. (2023). The S5, RWKV, and baseline results presented here are computed, following Nagy et al. (2023), for Alphabet (GOOG) and Intel (INTC) stock on a sub-sample of the test data from January 2023. The Coletta model is trained on three days from January 2019 and tested on three subsequent days, following the procedure in Coletta et al. (2022), which is necessary due to the high computational cost for training and inference with Coletta. Comparing all models, we observe that the LOBS5 model provides state-of-the-art performance on the benchmark task.
Figure 3 presents a key benchmark feature to compare multiple models across multiple score dimensions, allowing an examination of individual strengths and weaknesses. To provide summary scores per model, Figure 4 reports the mean, median, and inter-quartile mean for the L1 and Wasserstein-1 metrics for all models  [^2] The Coletta model (Coletta et al., 2022) was trained on both GOOG and INTC data, but failed to produce reasonable results for INTC, which is expected since the model was designed for small-tick stocks, whereas INTC is not.. Error bars demarcate the 99% bootstrapped confidence intervals. Metrics for individual scoring functions are shown in Figure 12 (Appendix E).
The benchmark also measures error divergence by comparing distributions of scoring functions, conditional on the inference time step. These demonstrate the rate at which distributions diverge from real data. Results show increasing errors across all models with the fastest divergence exhibited by the RWKV models across most scores. For the S5 model in particular, scoring functions with a dependence on features of the book states, which only gradually change, such as book volume, are expected to diverge, as the initial real data seed decays. However, the rate of decay can still be compared between models. See Figure 18 in appendix E for L1 divergence curves and appendix D for a discussion of and an ablation experiment on the effect of bin sizes on divergence scores.
@@
Figure 6: $\mathbf{L_{1}}$ distance between real and generated data histograms (incl. 99% CIs). baseline performs well on LOB depth and level-related scores, and much worse on time and volume metrics. LOBS5 dominates L1 loss for GOOG.
@@
Figure 7: Comparison of tick-normalized mid-price impact response functions for different event types between real and generated data-sets. Shaded regions are 99% confidence intervals. Compared are the LOBS5 and stochastic baseline models. In contrast to the baseline, LOBS5 reproduces most features of the expected impact response functions.
Figure 8: Comparison of F1-scores for the MLP classification models trained on historical (real) data only (solid lines), and trained on both real data and data generated (dashed lines) for a subset of the generated models. The error bars depict the 95% confidence intervals bootstrapped from five different training seeds.
@@
The response functions for Alphabet (GOOG) are shown in  Figure 7 for the Baseline and LOBS5 models. The LOBS5 model generally reproduces curves similar to real data but does so better for small-tick stock GOOG. In contrast, the baseline model (Cont et al., 2010) cannot faithfully reproduce impact curves. We do not post the impact curves for the Coletta and RWKV models, as they quickly diverge due to error accumulation in inference. The average $L_{1}$ distance between real and generated impact curves (see equation (7)) is $\Delta R=\textbf{2.45}$ for LOBS5, and $\Delta R=\textbf{126}$ for RWKV-6. For the LOBS5 model, we observe differences mainly in the $MO$ orders at short lags. This is due to the JAX-LOB simulator (Frey et al., 2023), which is used by the LOBS5 model during inference. JAX-LOB splits limit orders, which can only be partially filled, into execution messages and additional resting limit orders, thereby merging a multilevel midprice change into a single order book update.
F1-scores for mid-price trend forecasting using a simple MLP, following the LOBCAST implementation (Prata et al., 2024), are illustrated for Alphabet (GOOG) in Figure 8. We observe that, for the Coletta model in particular, including generated synthetic data markedly reduces the resulting prediction precision and recall. This is in line with model rankings based on the distributional distances shown in Figure 5. Generally, the same pattern is also present across the other models, particularly for short prediction horizons, although to a less extreme extent. Mixing generated data from either the baseline or LOB-S5 models into the training set has no significant effect for longer prediction horizons. Although it is desirable for models to generate data which increase prediction performance, not acting detrimentally is a minimum expectation. The results show that current generative models are not yet able to generate data aiding this mid-price trend prediction task. Beyond this conclusion, further quality differentiation between the models based on this task is limited. In contrast, the distributional LOB-Bench evaluation results provide a clearer picture of model strengths and potential for future improvement.
@@ 7 Conclusions
One of the interesting aspects of generative AI models for microstructure data is the ability to model counterfactuals, which is closely related to the notion of price impact in financial modelling. Factoring in the reactions of other market participants to one’s actions with conventional approaches is very challenging, but our benchmark suite for generative LOB models provides extensive tests to evaluate whether generated data reproduces the expected response functions at a larger scale. Future research could involve measuring the extent to which generative models match the market impact laws in the literature, such as the “square root law” (SRL) (Tóth et al., 2016). We hope that LOB-Bench opens the door to many new studies, including the training of reinforcement learning algorithms and multi-agent models for tasks such as trade execution, with the ability to model realistic reactions of different market participants.
@@ Acknowledgements
We gratefully acknowledge the Oxford-Man Institute of Quantitative Finance and the Foerster Lab for AI Research for providing us access to their GPU compute servers. These computational resources were instrumental in carrying out the experiments and analyses presented in this work. AC acknowledges funding from a UKRI AI World Leading Researcher Fellowship (grant EP/W002949/1. AC and JF acknowledge funding from a JPMC Faculty Research Award. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Benchmark Code
The benchmark code can be found on GitHub at https://github.com/peernagy/lob_bench, additional resources can be found on the project website https://lobbench.github.io/.
The benchmark suite provides a convenient API functionality to evaluate model data for a range of scoring functions and metrics. A specification of such functions and loss metrics can be defined in a configuration dictionary, which can then be passed to a function performing the unconditional and conditional model evaluation. Similarly, the benchmark provides functions to compute the market impact curves, along with a mean L1 score. A default configuration dictionary, specifying the scoring functions reported here, evaluated using L1 and Wasserstein-1 loss, is similarly provided for easy reproducibility.
@@ Appendix B LOBS5 Training Details
Starting from the model introduced by Nagy et al. (2023), we have scaled up the model size by adding additional S5 layers, with a resulting parameter count of approximately 35M (compared to originally 6.3M). The training consisted of 100 epochs of shuffled data sequences from the entire year of 2022, training with a total training budget of 30.4 L40 days (3.8 days across 8 GPUs). Adam (Kingma & Ba, 2014) was used as an optimizer with a cosine learning rate schedule. Losses on the test set over the course of training are displayed in Figure 9.
@@ Appendix C RWKV Training Details
Appendix D Sensitivity of Divergence Metrics to the Bin Size
We adopt a dynamic bin size determined by the Freedman–Diaconis (FD) rule, which is specifically designed to adapt to the underlying data distribution. As a result, we do not anticipate significant sensitivity to the choice of bin size. This choice is further supported by a theoretical convergence property: the FD rule minimizes the integrated mean squared error (IMSE) between the histogram and the true data distribution (Freedman & Diaconis, 1981).
To empirically evaluate the impact of different bin sizes on the reported divergence scores, we conduct the following experiments. The same divergence metric computations are repeated with half the regular bin size, as well as with double the regular bin size. To illustrate low bin size sensitivity, we perform this experiment for the LOBS5 model across all evaluated divergence metrics. We note that smaller bin sizes tend to increase both L1 and Wasserstein-1 errors, as data points, which would have been grouped into a single bin, could then be part of two separate bins. On the contrary, larger bin sizes tend to decrease errors, as data points close to each other are more likely to be grouped together.
To empirically assess the robustness of our results to bin size variation, we repeat the divergence metric computations using both half and double the default FD bin size. This analysis is conducted using the LOBS5 model across all evaluated divergence scores. We observe that reducing the bin size generally increases the $L_{1}$ errors. This is likely because data points that would have previously fallen within the same bin may now fall into separate bins, amplifying divergence estimates. Conversely, increasing the bin size tends to reduce these errors, as nearby data points are more likely to be aggregated into the same bin, leading to smoother approximations.
Figure 11 shows the mean divergence scores for the LOBS5 model, along with 99% confidence intervals. Doubling or halving the bin size results in a consistent decrease or increase in error scores, respectively. Notably, the magnitude of these deviations is comparable to the width of the confidence intervals under the default FD bin size. Given that these changes are relatively small and systematic, we conclude that while the choice of a theoretically grounded binning rule is important, model ranking remains stable as long as the same binning strategy is applied uniformly across all evaluated models.
Figure 11: L1 Divergence Scores with half and double regular bin size. Large bin size deviations (halfing or doubling) systematically affect the level of scores but not their ranking.
Appendix E Additional Figures
@@
Figure 15: rwkv4 - histograms comparing score distributions for real (blue) and generated (orange) LOB data for Alphabet (GOOG) and Intel (INTC) stocks. The model produces volatile data with larger spreads, missing correct order levels, leading to difficulty matching book volumes.